# Fusing Bluetooth with Pedestrian Dead Reckoning: A Floor Plan-Assisted Positioning Approach

### v2

@@ IV-A1 Stylization
To eliminate the distractions of engineering drawings, we redraw a stylized floor plan for positioning. In particular, the floor plan map features required by the algorithm often include Wall (i.e. non-passable obstacles) and Door (i.e. conditionally passable obstacles). We use distinguishable colors to fill the Wall and Door areas, such as black and red, while the Walkable areas are filled with white, as illustrated in Fig. 3(a).
@@ IV-A2 Feature Extraction
To identify the Rooms and obtain the precise areas of different map features, we need to extract these features from the floor plan. A mask-based method simplifies this process effectively. Adding the binary masks[42] of Walls, $\bm{I}_{\textit{Wall}}$, and Doors, $\bm{I}_{\textit{Door}}$, we obtain the mask of the whole obstacle area $\bm{I}_{\textit{obs}}$ by $\bm{I}_{\textit{obs}}=\bm{I}_{\textit{Wall}}+\bm{I}_{\textit{Door}}$. Then, we extract the inner contours of $\bm{I}_{\textit{obs}}$, which represent either a Room or an obstacle boundary within the Room, as shown in Fig. 3(b).
@@ IV-A4 Grid Array Construction
[…]
All of the above steps can be completed during the offline phase. Additionally, downsampling can be performed before stylization when the size of the floor plan is too large. This is because a large floor plan will lead to a large resolution factor $r$. Although this results in a higher map resolution, it also leads to larger computational costs. Therefore, it is necessary to downsample the floor plan, and set $r$ to an appropriate value without affecting the poisitioning accuarcy.
@@ IV-B Pedestrian Dead Reckoning (PDR)
[…]
@@ IV-C Fusion with PF
$p(\bm{x}^{(k)}|\mathcal{Z}^{(k)})\approx\textstyle\sum\limits_{i=1}^{m}\widetilde{w}_{i}^{(k)}\delta(\bm{x}^{(k)}-\widehat{\bm{x}}^{(k)}_{i}),$ (35a)
@@
where $\delta(\cdot)$ is the Dirac function. According to (5), the normalized particle weights $\{\widetilde{w}_{i}^{(k)}\}_{i=1}^{m}$ should be calculated as normalized $p(\bm{x}_{B}^{(k)}|\bm{x}_{i}^{(k)})$. Therefore, based on the assumption of Gaussian noise[39], these weights are calculated by:
@@
IV-D Floor Plan Assisted Post Processing
IV-D1 Correction Determination
We first define the Raycast function, which is summarized in Function 1. In Raycast, we move from a start point incrementally towards the end point, until it is reached. If an obstacle is encountered during this process, the Raycast returns $\mathbf{True}$ and the first contact point $\bm{h}$ (also referred to as the hit point). Raycast efficiently detects whether there is an obstacle between two points; if an obstacle exists, it can find the hit point with a time complexity of $\mathcal{O}(n)$.
@@ IV-D1 Correction Determination
         [Update step: $\delta=0.05$];
@@ 1 $\texttt{hit}=\mathbf{False}$;
9             return $\texttt{hit}=\mathbf{True}$, $\bm{h}=\bm{x}_{1}+k\delta\bm{x}_{\mathrm{n}}$;
@@
12if !hit then return $\texttt{hit}=\mathbf{False}$, $\bm{h}=\mathbf{Null}$;
@@
At a time $k$, perform Raycast from the previous position $\bm{x}^{(k-1)}$, to the estimated position $\widehat{\bm{x}}$. If the hit point $\bm{h}$ is obtained, it is considered that moving from $\bm{x}^{(k-1)}$ to $\widehat{\bm{x}}$ is not feasible. The infeasible step necessitates a position correction, which typically includes step length correction and yaw correction. Then, a candidate rotation angle sequence $\mathbf{\Phi}=\{\pm i\phi\}_{i=1}^{N}$ is defined. In this sequence, $\phi$ is a unit rotation angle, and the elements in $\mathbf{\Phi}$ are arranged in ascending order of their absolute values.
@@ IV-D2 Position Correction
(i) If the yaw correction is needed, as shown in Fig. 5, we find the optimal correction rotation angle by sequentially traversing all candidate angles in $\mathbf{\Phi}$ and then determining the optimal corrected position. $\forall\phi^{\prime}\in\mathbf{\Phi}$, we can construct a rotation matrix by:
$\bm{R}(\phi^{\prime})\triangleq\begin{bmatrix}\cos\phi^{\prime}&-\sin\phi^{\prime}\\ \sin\phi^{\prime}&\cos\phi^{\prime}\end{bmatrix},$ (38)
and rotate $\widehat{\bm{s}}$ to obtain a test point $\bm{x}_{\mathrm{t}}(\phi^{\prime})$:
$\bm{x}_{\mathrm{t}}(\phi^{\prime})=\bm{x}^{(k-1)}+\bm{R}(\phi^{\prime})\cdot f\widehat{\bm{s}},$ (39)
where $f\geq 1$ is a scale factor. Next, we follow the order of $\mathbf{\Phi}$, sequentially searching for the optimal rotation angle $\phi^{*}$ with the smallest absolute value, such that Raycast from $\bm{x}^{(k-1)}$ to $\bm{x}_{\mathrm{t}}(\phi^{\prime})$ returns $\mathbf{False}$, which indicates that this path is feasible without any obstruction:
$\phi^{*}$ $=\arg\min_{\phi^{\prime}\in\mathbf{\Phi}}\,|\phi^{\prime}|$ (40)
$\mathrm{s.t.}~{}~{}\mathrm{Raycast}$ $(\bm{M},\bm{x}^{(k-1)},\bm{x}_{\mathrm{t}}(\phi^{\prime}))=\mathbf{False}.$ (40a)
@@
$\bm{x}^{(k)}=\bm{x}^{(k-1)}+\bm{R}(\phi^{*})\cdot\widehat{\bm{s}}.$ (41)
If a suitable rotation angle cannot be found within sequence $\mathbf{\Phi}$, i.e., $\forall\phi^{\prime}\in\mathbf{\Phi},~{}\mathrm{Raycast}(\bm{M},\bm{x}^{(k-1)},\bm{x}_{\mathrm{t}}(\phi^{\prime}))=\mathbf{True}$, we can only assume that the step is blocked by an obstacle. As shown in Fig. 5, in this case, a step length correction is applied instead, by moving a small distance backward from the hit point $\bm{h}$ to Walkable area, yielding the corrected position:
@@
1 Set candidate angles as $\mathbf{\Phi}=[\pm j\phi~{}\textbf{for}~{}j=1,2,\dots,N]$;
2 Obtain flag hit and point $\bm{h}$ by $\mathrm{Raycast}(\bm{M},\bm{x}^{(k-1)},\widehat{\bm{x}})$;
@@ 8             Obtain the incident angle $\alpha$ by (37);
10                   $\alpha_{+}=\langle\bm{n}_{h},-\bm{R}(\phi_{+})\cdot\widehat{\bm{s}}\rangle$;
11                   $\alpha_{-}=\langle\bm{n}_{h},-\bm{R}(\phi_{-})\cdot\widehat{\bm{s}}\rangle$;
@@ 32      Update the particles by (44);
Since all the internal contours of the map are obtained in the offline phase in Section IV-A2, it is easy to determine the Room of the last result $\bm{x}^{(k-1)}$, i.e. the contour that contains $\bm{x}^{(k-1)}$. Since the estimated results of GML algorithm are typically accurate enough to distinguish different Rooms, we use $\bm{x}_{B}$ as the basis for determining the current Room. As shown in Fig. 6, if $\bm{x}_{B}$ is still in the previous Room, we still consider it to be blocked by the Door, and thus, the result is corrected by moving from the hit point backward to the Walkable area. Otherwise, in Fig. 6, we consider the user to pass through the Door in this step. We identify the closest point $\bm{x}_{\mathrm{c}}$ from all the contours, and correct the result by moving a small step forward from $\bm{x}_{\mathrm{c}}$ into Walkable area, completing the action of passing through the Door:
@@ V Experimental Setup and Results
TABLE I: Key Experimental Parameters
Device Parameter Value Algorithm Parameter Value
BLE Beacon Model RF-star RF-B-SR1 FP-BP Floor plan resolution $r=1/0.07~{}\mathrm{m^{-1}}$
Broadcast interval $200~{}\mathrm{ms}$ Virtual grid interval $I_{M}=0.3~{}\mathrm{m}$
Transmission power $0~{}\mathrm{dBm}$ BLE estimation interval $250~{}\mathrm{ms}$
Advertising mode iBeacon BLE Kalman filter $q_{K}=0.16$, $r_{K}=16$
Smartphone Model HONOR Magic 6 PF particle number $m=500$
Mobile processor Snapdragon 8 Gen3 PF particle noise $|\nu_{s}|\leq 0.05~{}\mathrm{m}$, $|\nu_{\theta}|\leq 10^{\circ}$
Operating system Android 15 PF fusion variance $\sigma_{x}^{2}=0.1$
BLE scan mode Low latency PPC candidate angles $\mathbf{\Phi}=\{\pm i\cdot 5^{\circ}\}_{i=1}^{12}$
IMU sample rate $60~{}\mathrm{Hz}$ PPC test scale factor $f=1.5$
Figure 8: a) BLE beacons in the experimental area; b) FP-BP prototype on the smartphone.
In this section, a prototype is established to implement FP-BP, and experiments are conducted to verify its performance and feasibility. Fig. 8 shows the experimental area and prototype of FP-BP, and key experimental parameters are listed in Table I. Next, we will detail the implementation of the experiments and analyze the results.
@@ V-A Experimental Setup
The experimental area is located in a building with a ceiling height of approximately $2.5~{}\mathrm{m}$, and a total area of about $5760~{}\mathrm{m}^{2}$. In our experiments, a total of $39$ beacons are deployed on the ceiling of this floor. These beacons, as shown in Fig. 8, have distinct UUIDs and are deployed at the coordinates shown in Fig. 9.
Figure 9: Beacon deployment in the experimental area.
Figure 10: A brief diagram of the experimental trajectories.
In the experiments, we set the tester to walk along a specific trajectory at a speed of approximately $1.0~{}\mathrm{m/s}$, holding the smartphone facing forward (i.e., the $y^{\mathrm{D}}$-axis of the phone is directed forward), with the step length fixed at $0.6~{}\mathrm{m}$. To verify the robustness of FP-BP, we set up three trajectories, as shown in Fig. 10. The total lengths of these trajectories are $316.8~{}\mathrm{m}$, $137.4~{}\mathrm{m}$ and $187.8~{}\mathrm{m}$, respectively. In particular, for trajectory 1, the tester walks two laps for one trial; for trajectory 2, there are two Room switch events, whereas other trajectories have none. Among them, trajectory 1 is considered the most representative, and thus Sections V-B and V-C present the results of $10$ trials of trajectory 1 conducted by two testers. The comparisons across these three trajectories are presented in Section V-D.
To evaluate the positioning performance, we use the following metrics: (i) The 2-D position error (PE), which is defined as the distance between the estimated position and the ground truth in MCS. (ii) The average of multiple PEs, termed MPE. (iii) The average of multiple computation times, termed MCT.
Figure 11: Experimental results of trajectory 1. (a)–(c): Comparison of GML with BLE baseline algorithms. (d)–(f): Comparison of FP-BP in different cases. (g)–(i): Comparison of FP-BP with baseline algorithms.

### v3

@@ IV-A1 Stylization
To eliminate the distractions of computer-aided design (CAD) drawings, we manually redraw a stylized floor plan for positioning. In particular, the floor plan map features required by the algorithm often include Wall (i.e., non-passable obstacles) and Door (i.e., conditionally passable obstacles). We use distinguishable colors to fill the Wall and Door areas, such as black and red, while the Walkable areas are filled with white, as illustrated in Fig. 33(a).
@@ IV-A2 Feature Extraction
To identify the Rooms and obtain the precise areas of different map features, we need to extract these features from the floor plan. A mask-based method simplifies this process effectively. Adding the binary masks[43] of Walls, $\bm{I}_{\textit{Wall}}$, and Doors, $\bm{I}_{\textit{Door}}$, we obtain the mask of the whole obstacle area $\bm{I}_{\textit{obs}}$ by $\bm{I}_{\textit{obs}}=\bm{I}_{\textit{Wall}}+\bm{I}_{\textit{Door}}$. Then, we extract the inner contours of $\bm{I}_{\textit{obs}}$, which represent either a Room or an obstacle boundary within the Room, as shown in Fig. 33(b). Note that this step simply determines the map feature of each pixel individually based on its specific color, which is independent of the overall feature shape, and thus can avoid misjudgment.
@@ IV-A4 Grid Array Construction
In Section III, we have proposed the GML algorithm, which selects the grid point with the maximum likelihood probability as the positioning result. To apply the GML algorithm for BLE positioning, we need to construct a virtual grid array in advance, where the grid points are distributed over the entire map. Also, we should remove all the grid points within obstacle areas. Let the map range be $0,X]\times[0,Y]$, and the grid interval be $I_{M}$. Then, the position of the $(i,j)$-th grid point is $\bm{g}_{ij}=[(i+0.5)X/I_{M},(j+0.5)Y/I_{M}]^{\mathsf{T}}$, where $1\leq i\leq\lfloor X/I_{M}\rfloor-1$, $1\leq j\leq\lfloor Y/I_{M}\rfloor-1$, and $\lfloor\cdot\rfloor$ denotes the floor function. Let the grid array be denoted as $\mathcal{G}$, and then all grid points in $\mathcal{G}$ are located in Walkable area, i.e., $\mathcal{G}=\{\bm{g}_{ij}\mid\bm{M}(\bm{g}_{ij})=0\}$. Substituting this grid array $\mathcal{G}$ into (22) and (23), we can ensure that the BLE estimation result is located solely within Walkable area.
All of the above steps can be completed during the offline phase. Additionally, we note that: (i) When the floor plan is too large, downsampling can be applied before Stylization to avoid an excessively high map resolution factor $r$, which would otherwise increase computational cost. (ii) If the building structure changes, only Stylization requires manual adjustment, while the remaining steps are fully automated, thereby minimizing system update and maintenance costs.
@@ IV-B Pedestrian Dead Reckoning (PDR)
where $h$ is half of the window size; $z_{\mathrm{th}}$ is the threshold for maximum step acceleration; $k_{\mathrm{th}}$ is the threshold for minimum step interval; and $h^{\prime}$ is the last step event time. Only when these three conditions are simultaneously satisfied can time khk-h be recognized as a step event. At this time, the peak-to-peak value of the acceleration within this $2h$ window is calculated as:
@@ IV-C Fusion with PF
$p(\bm{x}^{(k)}|\mathcal{Z}^{(k)})\approx\textstyle\sum\limits_{i=1}^{m}\widetilde{w}_{i}^{(k)}\delta_{D}(\bm{x}^{(k)}-\widehat{\bm{x}}^{(k)}_{i})$, (37a)
@@
where $\delta_{D}(\cdot)$ is the Dirac function. According to (5), the normalized particle weights $\{\widetilde{w}_{i}^{(k)}\}_{i=1}^{m}$ should be calculated as normalized $p(\bm{x}_{B}^{(k)}|\bm{x}_{i}^{(k)})$. Therefore, based on the assumption of Gaussian noise[39], these weights are calculated by:
@@
   [Update step: $\delta<r^{-1}$];
@@ 1 $\texttt{hit}=\mathbb{F}$;
9     return $\texttt{hit}=\mathbb{T}$, $\bm{h}=\bm{x}_{1}+k\delta\bm{x}_{\mathrm{n}}$;
@@
12if !hit then return $\texttt{hit}=\mathbb{F}$, $\bm{h}=\mathbf{Null}$;
@@
IV-D Floor Plan Assisted Post Processing
IV-D1 Correction Determination
We first define the Raycast function, which is summarized in Function 1. In Raycast, we move from a start point incrementally towards the end point, until it is reached. If an obstacle is encountered during this process, the Raycast returns $\mathbb{T}$ and the first contact point $\bm{h}$ (also referred to as the hit point). Raycast efficiently detects whether there is an obstacle between two points; if an obstacle exists, it can find the hit point with a time complexity of $\mathcal{O}(n)$. Note that the update step $\delta$ should be set such that $\delta<r^{-1}$ to prevent missed detections, where $r$ is the map resolution, since any obstacle thinner than one pixel (i.e., $r^{-1}$ in reality) would otherwise be ignored in the image.
At a time $k$, perform Raycast from the previous position $\bm{x}^{(k-1)}$, to the estimated position $\widehat{\bm{x}}$. If the hit point $\bm{h}$ is obtained, it is considered that moving from $\bm{x}^{(k-1)}$ to $\widehat{\bm{x}}$ is not feasible. The infeasible step necessitates a position correction, which typically includes step length correction and yaw correction. Then, a candidate rotation angle sequence $\mathbf{\Phi}=\{\pm i\cdot\Delta\phi\}_{i=1}^{N}$ is defined. In this sequence, $\Delta\phi$ is a unit rotation angle, and the elements in $\mathbf{\Phi}$ are arranged in ascending order of their absolute values.
@@ IV-D2 Position Correction
(i) If the yaw correction is needed, as shown in Fig. 55(a), we find the optimal correction rotation angle by sequentially traversing all candidate angles in $\mathbf{\Phi}$ and then determining the optimal corrected position. $\forall\phi\in\mathbf{\Phi}$, we can construct a rotation matrix by:
$\bm{R}_{\phi}\triangleq\begin{bmatrix}\cos\phi&-\sin\phi\\ \sin\phi&\cos\phi\end{bmatrix}$, (40)
and rotate $\widehat{\bm{s}}$ to obtain a test point $\bm{t}_{\phi}$:
$\bm{t}_{\phi}=\bm{x}^{(k-1)}+\bm{R}_{\phi}\cdot f\widehat{\bm{s}}$, (41)
where $f\geq 1$ is a scale factor. Next, we follow the order of $\mathbf{\Phi}$, sequentially searching for the optimal rotation angle $\phi^{*}$ with the smallest absolute value, such that Raycast from $\bm{x}^{(k-1)}$ to $\bm{t}_{\phi}$ returns $\mathbb{F}$, which indicates that this path is feasible without any obstruction:
$\phi^{*}=\arg\min_{\phi\in\mathbf{\Phi}}\,|\phi|$ (42)
$\mathrm{s.t.}~~\mathrm{Raycast}(\bm{M},\bm{x}^{(k-1)},\bm{t}_{\phi})=\mathbb{F}.$ (42a)
@@
$\bm{x}^{(k)}=\bm{x}^{(k-1)}+\bm{R}_{\phi^{*}}\cdot\widehat{\bm{s}}$. (43)
If a suitable rotation angle cannot be found within sequence $\mathbf{\Phi}$, i.e., $\forall\phi\in\mathbf{\Phi},~\mathrm{Raycast}(\bm{M},\bm{x}^{(k-1)},\bm{t}_{\phi})=\mathbb{T}$, we can only assume that the step is blocked by an obstacle. As shown in Fig. 55(b), in this case, a step length correction is applied instead, by moving a small distance backward from the hit point $\bm{h}$ to Walkable area, yielding the corrected position:
@@
1 Set candidate angles as $\mathbf{\Phi}=[\pm j\cdot\Delta\phi~\textbf{for}~j=1,2,\dots,N]$;
[…]
@@ 8     Obtain the incident angle $\alpha$ by (39);
10        $\alpha_{+}=\langle\bm{n}_{h},-\bm{R}_{\phi_{+}}\cdot\widehat{\bm{s}}\rangle$;
11        $\alpha_{-}=\langle\bm{n}_{h},-\bm{R}_{\phi_{-}}\cdot\widehat{\bm{s}}\rangle$;
@@ 32 Update the particles by (46);
Since all the internal contours of the map are obtained in the offline phase in Section IV-A2, it is easy to determine the Room of the last result $\bm{x}^{(k-1)}$, i.e., the contour that contains $\bm{x}^{(k-1)}$. Since the estimated results of GML algorithm are typically accurate enough to distinguish different Rooms, we use $\bm{x}_{B}$ as the basis for determining the current Room. As shown in Fig. 66(a), if $\bm{x}_{B}$ is still in the previous Room, we still consider it to be blocked by the Door, and thus, the result is corrected by moving from the hit point backward to the Walkable area. Otherwise, in Fig. 66(b), we consider the user to pass through the Door in this step. We identify the closest point $\bm{x}_{\mathrm{c}}$ from all the contours, and correct the result by moving a small step forward from $\bm{x}_{\mathrm{c}}$ into Walkable area, completing the action of passing through the Door:
@@
TABLE I: Key Experimental Parameters
Device Parameter Value FP-BP Algorithm Parameter Value
BLE beacon Model RF-star RF-B-SR1 Floor plan Resolution $r=1/0.07~\mathrm{m^{-1}}$
Broadcast interval $200\,\mathrm{ms}$ Virtual grid interval $I_{M}=0.3\,\mathrm{m}$
Transmission power $0\,\mathrm{dBm}$ GML Estimation interval $250~\mathrm{ms}$
Advertising mode iBeacon Kalman filter $q_{K}=0.16$, $r_{K}=16$
Smartphone Model HONOR Magic 6 Tunable parameters $\kappa=0.01$, $\tau=0.5$
Mobile processor Snapdragon 8 Gen3 PF fusion Particle number $m=500$
Operating system Android 15 Particle noise $|\nu_{s}|\leqslant 0.05\,\mathrm{m}$, $|\nu_{\theta}|\leqslant 10^{\circ}$
RAM 12 GB Fusion variance $\sigma_{x}^{2}=0.1$
BLE scan mode Low latency PPC Candidate angles $\mathbf{\Phi}=\{\pm i\cdot 5^{\circ}\}$, $1\leqslant i\leqslant 12$
IMU sample rate $60\,\mathrm{Hz}$ Test scale factor $f=1.5$
@@
IV-E Extending FP-BP for Multi-Floor Scenarios
The above algorithm is designed for single-floor scenarios. In multi-floor positioning, users often move across floors via elevators or stairwells. We note that the proposed FP-BP can be readily extended to such scenarios.
First, we can mark areas involving floor transitions (FTAs) in the floor plan, e.g., elevators and stairwells, with a new distinguishable color during the preprocessing step. This makes it straightforward to extract their binary mask $\bm{I}_{\textit{FTA}}$. Next, we can extend the binary mask $\bm{I}(\bm{x}^{\mathrm{I}})$ in (26) by defining $\bm{I}=\bm{I}_{\textit{Wall}}+0.5\bm{I}_{\textit{Door}}+0.25\bm{I}_{\textit{FTA}}$, and the map $\bm{M}(\bm{x})$ in (27) by defining $\bm{M}(\bm{x})=0.25$ if $\bm{x}$ is in the FTA. When the estimated position $\bm{x}^{(k)}$ enters the FTA, a floor determination is performed at each subsequent update, primarily based on BLE signals. Owing to their significant cross-floor attenuation[4], proximity to multiple beacons on another floor indicates a floor transition. In addition, if the smartphone is equipped with a barometer, the altitude can be estimated to further assist in floor determination[48].
Once the user is detected to have entered another floor, the system switches to the corresponding floor plan and re-initializes the algorithm, including resetting the initial position based on BLE and re-initializing the PF, so that a new round of single-floor positioning can be performed.
Figure 8: (a) BLE beacons in the experimental area; and (b) FP-BP prototype on the smartphone.
Figure 9: BLE beacons in the experimental area. Only the beacons with “*” are considered in the experiments under the sparse beacon deployment.
@@ V Experimental Setup and Results
In this section, a prototype is established to implement FP-BP, and experiments are conducted to verify its performance and feasibility. Fig. 8 shows the experimental area and prototype of FP-BP, and key experimental parameters are listed in Table I unless otherwise specified. Next, we will detail the implementation of the experiments and analyze the results.
@@ V-A Experimental Setup
The experimental area was located on the 6th and 7th floors of a teaching building, with a ceiling height of approximately $2.5\,\mathrm{m}$ and a total area of about $5760\,\mathrm{m}^{2}$ per floor. As shown in Fig. 88(a), in our experiments, a total of $56$ beacons are deployed on the ceiling, which have distinct UUIDs and are deployed at the coordinates shown in Fig. 9.
(a) Trajectories 1–3.
(b) Cross-floor trajectory 4.
Figure 10: Brief diagrams of experimental trajectories.
In the experiments, we set the tester to walk along a specific trajectory at a speed of approximately $1.0~\mathrm{m/s}$, holding the smartphone facing forward (i.e., the $y^{\mathrm{D}}$-axis of the phone is directed forward), with the step length fixed at $0.6\,\mathrm{m}$. To verify the robustness of FP-BP, we set up four trajectories, as shown in Fig. 10. The total lengths of these trajectories are $316.8\,\mathrm{m}$, $137.4\,\mathrm{m}$, $187.8\,\mathrm{m}$ and $115.8\,\mathrm{m}$, respectively. In particular, for trajectory 1, the tester walked two laps for one trial; for trajectory 2, there are two Room switch events, whereas other trajectories have none; and for trajectory 4, the tester started on the 6th floor and was transported by elevator to the 7th floor. The following Sections V-B and V-C present the results of $10$ trials of trajectory 1 conducted by two testers; Section V-D presents the results and comparisons across these trajectories; and finally, Section V-E compares the performance of FP-BP on different mobile devices.
To evaluate the positioning performance, we first define the 2-D position error as the distance between the estimated position and the ground truth in MCS; then use the following metrics: (i) The mean of errors (MPE); (ii) 50%/80% cumulative distribution function (CDF) corresponding errors (P50/P80 scores); (iii) the standard deviation (STD) of errors; and (iv) the mean of multiple computation times (MCT).
Figure 11: Experimental results of trajectory 1. (a) and (b): Comparison of GML with BLE baseline algorithms. (c) and (d): Ablation comparison of FP-BP in different cases. (e) and (f): Comparison of FP-BP with baseline algorithms.